OpenStudy (anonymous):

A tree casts a shadow of 100 inches, while a 7-inch branch casts a shadow of 10 inches. Using similar triangles, determine how tall the tree is.

OpenStudy (nurali):

A tree casts a shadow of 100 inches while a 7 inch branch casts a shadow of 10 inches, how tall is the tree? shadow 1 / shadow 2 = height 1 / height 2 100 / 7 = x / 10 1000 / 7 = x (multiplied both sides by 10) 142.857 inches approx = x

OpenStudy (anonymous):